Consider the following data for question 1-6: Two solutions of an unknown slightly soluble salt, A(OH)2, were allowed to equilibrate -- one at 25oC and the other at 80 oC. A 15.00 mL aliquot of each solution is titrated with 0.200 M HCl. 9.00 mL of the acid is required to reach the endpoint of the titration at 25 oC, while 74.35 mL are required for the 80 oC solution.

Calculate the Ksp at 25 oC.
Answer #1

Answer:

To calculate the Ksp at 25

Reaction is given as follows

HCl moles = molarity * volume

= 0.200 * 9

= 1.8 mmol

Here it is given that 15 ml aliquot is titrated with HCl

2X * 15 = 1.8 mmol

X = 0.06M

Now ,

Ksp = [A+][OH-]^2

= x*(2x)^2

= 4x^3

= 4(0.06)^3

Ksp = 0.000864